40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 541     Entry time: Wed Jun 18 18:26:19 2008
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: Finding the optimal operation temperature for the NPRO by the slow act scan 
Being suspicious of the temperature stabilization of the NPRO crystal, I ran the slow scan script written by Rana to find the suitable operation temperature.
The procedure is the same as the one explained in the entry below:
http://www.ldas-sw.ligo.caltech.edu/ilog/pub/ilog.cgi?group=40m&task=view&date_to_view=09/04/2006&anchor_to_scroll_to=2006:09:04:22:23:56-rana
The attached plots show the results. By looking at C1:PSL-126MOPA_126MON, I set the slow slider voltage to 0.
This time, it looks like the temperature control of the NPRO crystal is working fine.
Obviously, PMC picks up many higher order modes. I will try to mode match/align the PMC later.
Attachment 1: FSS-slow-scan.pdf  229 kB  | Hide | Hide all
FSS-slow-scan.pdf
ELOG V3.1.3-